Please Help!! Due Today, need to know how to solve it, this time cant even figure out the table, How many kilograms of zinc that is 18% pure should be mixed wth 16 kilograms of alloy that is 24% pure to make an alloy that is 22% pure zinc?
Found 2 solutions by Alan3354, nerdybill:
Answer by Alan3354(31538)

How many kilograms of zinc that is 18% pure should be mixed wth 16 kilograms of alloy that is 24% pure to make an alloy that is 22% pure zinc?
--------------------
Add x kg of 18%, = 0.18x
Added...Starting....Result
0.18x + 16*0.24 = (x+16)*0.22
-----------------------------
x100 to get rid of fractions
18x + 384 = 22x + 352
-4x = -32
x = 8 kg of 18%

Answer by nerdybill(7012)

How many kilograms of zinc that is 18% pure should be mixed wth 16 kilograms of alloy that is 24% pure to make an alloy that is 22% pure zinc?
.
Let x = kilograms of zinc
then
.18x + .24(16) = .22(x+16)
.18x + 3.84 = .22x+3.52
0.32 = .04x
8 kilograms = x